OpenStudy (luigi0210):

Evaluate the integral:

OpenStudy (nincompoop):

0

OpenStudy (luigi0210):

\[\LARGE \int_{0}^{\pi/2}~sin^2x~cos^2x~dx\]

OpenStudy (nincompoop):

try without the limits first

OpenStudy (anonymous):

I'll do the indefinite you do the limits :D

OpenStudy (nincompoop):

copy cat

OpenStudy (nincompoop):

copy bat*

OpenStudy (anonymous):

Actually no, luigi show me what you've done first...

OpenStudy (anonymous):

Can you do it by parts?

OpenStudy (nincompoop):

you can do anything you want

OpenStudy (anonymous):

cos^2x = (1-sin^2x)

OpenStudy (anonymous):

\[\int\limits \sin^2x(1-\sin^2x)dx => \int\limits (\sin^2x - \sin^4x)dx \]

OpenStudy (anonymous):

This is the identity: \[\frac{ 1 }{ 4 }\sin^{2}(2x)\]

OpenStudy (anonymous):

Lets confuse luigi more

OpenStudy (luigi0210):

I changed it to \[\LARGE \int (\frac{1+cos2x}{2})(\frac{1-cos2x}{2})~dx\] but got the wrong answer..

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

that ultimately reduces to (1/4) sin^2(2x)

OpenStudy (anonymous):

\[\int\limits \sin^2xcos^2x dx = \int\limits \sin^2x(1-\sin^2x)dx\] Because cos^2x = 1-sin^2x, then you can expand it and get, \[\int\limits (\sin^2x-\sin^4x) dx = \int\limits \sin^2xdx - \int\limits \sin^4x dx\] You can use the reduction formula here, do you know it?

OpenStudy (kainui):

\[I=\int\limits_0^{\pi/2} \sin^2 x \cos^2 x\ dx\]\[J=\int\limits_0^{\pi/2} \sin^2 x \ dx\]\[J-I=\int\limits_0^{\pi/2} \sin^2 x(1-\cos^2 x) \ dx=\int\limits_0^{\pi/2} \sin^4 x \ dx\] ok now you can solve this more complicated simplification. GL!

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\sin^{2}(2x)=\frac{ 1 }{ 8 }(1-\cos(4 x)) \]

OpenStudy (nincompoop):

neat trick, kai

OpenStudy (kainui):

\[I=\int\limits_0^{\pi/2} \sin^2 x-\sin^4 x \ dx\] Although honestly is that any better?

OpenStudy (anonymous):

Don't bully me

OpenStudy (nincompoop):

laughing out loud he's batman

OpenStudy (kainui):

I'm robbin'

OpenStudy (nincompoop):

you can use power reduction from the get go

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\sin^3xcosx+\int\limits \sin^2xdx-\frac{ 3 }{ 4 } \int\limits \left( \frac{ 1 }{ 2 }-\frac{ 1 }{ 2 }\cos(2x) \right)dx\] eek

OpenStudy (kainui):

ok the real way to do it is like this:

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/2}\frac{ 1 }{ 4 }\sin^{2}(2x)=\int\limits_{0}^{\Pi/2}\frac{ 1 }{ 8 }(1-\cos(4x))\]

OpenStudy (kainui):

\[\int\limits_0^{\pi/2} \sin^2 x \cos^2 x \ dx=\int\limits_0^{\pi/2} (\sin x \cos x )^2\ dx=\int\limits_0^{\pi/2} (\frac{1}{2}\sin (2x) )^2 \ dx\]

OpenStudy (kainui):

\[\frac{1}{4}\int\limits_0^{\pi/2} \sin ^2 (2x) \ dx =\frac{1}{4}\int\limits_0^{\pi/2} \frac{1-\cos (4x)}{2} \ dx=\frac{1}{8}\int\limits_0^{\pi/2} 1-\cos (4x) \ dx\]

OpenStudy (anonymous):

\[\int\limits \sin^2xcos^2x dx = \frac{ 1 }{ 32 }(4x-\sin(4x))+C \] Final answer of indefinite integral, fyi.

OpenStudy (anonymous):

So pi/16 with limits

OpenStudy (kainui):

Actually there's more cleverness I suggest you do while solving this: 4x=u to get rid of how frequent it is. \[\frac{1}{2}\int\limits_0^{2\pi} 1-\cos (u)\ du\] Notice how that limit of integration changed? That automatically allows us to drop the cosine part! Why? Well every 2pi it repeats right? Just look at this  those areas cancel out every 2pi so don't even bother integrating it since you know it's garbage. \[\frac{1}{2}\int\limits_0^{2\pi}du=\pi\] That's right right?